### v1

# Disentangled Clothed Avatar Generation from Text Descriptions

Abstract
In this paper, we introduced a novel text-to-avatar generation method that separately generates the human body and the clothes and allows high-quality animation on the generated avatar. While recent advancements in text-to-avatar generation have yielded diverse human avatars from text prompts, these methods typically combine all elements—clothes, hair, and body—into a single 3D representation. Such an entangled approach poses challenges for downstream tasks like editing or animation. To overcome these limitations, we propose a novel disentangled 3D avatar representation named Sequentially Offset-SMPL (SO-SMPL), building upon the SMPL model. SO-SMPL represents the human body and clothes with two separate meshes, but associates them with offsets to ensure the physical alignment between the body and the clothes. Then, we design an Score Distillation Sampling(SDS)-based distillation framework to generate the proposed SO-SMPL representation from text prompts. In comparison with existing text-to-avatar methods, our approach not only achieves higher exture and geometry quality and better semantic alignment with text prompts, but also significantly improves the visual quality of character animation, virtual try-on, and avatar editing. Our project page is at https://shanemankiw.github.io/SO-SMPL/
Figure 1: Our method generates high-quality separated human body and clothes meshes from text prompts. Kinematics or simulation motions can drive the disentangled avatar representations to achieve photorealistic animations.
1 Introduction
Human avatars play a significant role in visual storytelling, conveying narratives through actions, expressions, and interactions with other elements in the animated world. This helps in creating a more dynamic and engaging storytelling experience for the audience. Avatars can be designed to represent diverse characters, including different genders, races, ethnicities, and abilities, making the animation more inclusive and accessible to a wider audience. This can also promote positive representation and challenge stereotypes in the media. However, the creation of high-quality avatars is typically a costly and labor-intensive endeavor, demanding the expertise of skilled 3D modelers.
In order to achieve a photorealistic, animatable, and customizable avatar, it is imperative to employ a disentangled representation consisting of distinct components, such as the body and clothing elements. Utilizing a disentangled human mesh facilitates enhanced customization and reusability of the avatar, enabling seamless alterations to clothing while maintaining the same underlying human body structure. Furthermore, the separate components contribute to heightened realism in animation, as different avatar parts exhibit distinct motion characteristics. For instance, human body movements can be accurately animated using forward kinematics, while cloth motion may be generated through physical simulations.
Despite the evident benefits, the majority of existing research primarily concentrates on reconstructing a singular avatar mesh that encompasses all components, derived from videos or images, consequently neglecting the potential for a disentangled representation. Only a limited number of studies, such as DELTA [26] and SCARF [25], have investigated the reconstruction of a disentangled avatar representation from video sources. However, these approaches aim to reconstruct a disentangled mesh but not facilitate the arbitrary creation of a disentangled avatar representation. As a result, the development of a comprehensive method for creating disentangled human avatars remains an unresolved challenge within the field.
Extending the existing text-to-avatar pipelines to generate disentangled human avatars is not trivial. The primary obstacle lies in not only independently generating the body and clothing components but also seamlessly combining them in a physically accurate manner while maintaining visually appealing results. Directly generating the body and clothing separately often results in misaligned components, detracting from the overall avatar quality. To overcome this issue, we introduce a novel representation called Sequentially-Offset-SMPL (SO-SMPL). SO-SMPL is based on the SMPL-X [53] as it effectively characterizes coarse human shapes within its parameter space. Then, SO-SMPL adds additional trainable offsets and appearance features to the SMPL-X model to represent the unclothed human body, similar to TADA [47]. Subsequently, to represent the clothing on the body, we build upon the human body to learn additional offsets, and appearances. This sequential representation ensures that the clothing is perfectly aligned with the human body while maintaining separate mesh encodings for each component.
Leveraging the power of stable diffusion with SDS losses [62], we introduce a novel approach to generate our SO-SMPL representation from text prompts. Our method achieves disentangled avatar generation by first creating the unclothed human body and subsequently distilling the clothing components based on the underlying body structure. Experiments show that our method produces high-quality human avatars with better geometry details and appearances than baseline methods [47, 32]. The disentangled nature of our SO-SMPL representation unlocks strong customization capabilities, allowing for the seamless integration of various clothing options on the same human body. Furthermore, the avatars generated by our state-of-the-art method exhibit exceptionally realistic animation results, achieved by simulating distinct motion characteristics for both clothing and human body components.
2 Related work
Distilling diffusion for 3D tasks. Many recent works have utilized diffusion models to generate 3D objects or scenes. Some works [80, 6, 40, 57, 63, 42, 51] trained 3D diffusion models to directly generate 3D representations such as point clouds, meshes, radiance fields[56], SDFs [79], DMTets[70] or tri-planes [11]. Others designed diffusion models to generate 3D-aware 2D images or textures[72, 50, 52, 49, 78, 66, 8, 28, 55, 12]. Both lines of works rely on 3D data[20, 19, 24] or multi-view images[65, 87, 91] to train.
With the success of 2D text-to-image diffusion models [30, 67, 21], recent works [62, 77, 5, 16, 48, 82, 68, 69, 13, 75, 36, 83, 85, 89, 101] propose to distill 3D representations from 2D diffusion models to circumvent the lack of 3D data. Some of them designed different 3D representations [48, 13, 75] for generation, while others explored the guidance schemes of diffusion models [62, 77, 82].
3D Human Avatar generation. Previous works on 3D avatar generation [14, 58, 27, 96, 31, 59, 7, 95] could generate diverse human geometries or textured avatars, but the generation process could not be controlled by text descriptions. With the help of large language models like CLIP [64] and text-to-image diffusion models [30, 67], recent works [9, 35, 99, 88, 43] were able to generate high-fidelity avatars with prompts or with both image and prompt [84, 38]. Other works have attempted to generate human avatars by painting the given human meshes with generated textures [73, 90]. However, the generated avatars could not be animated and utilized in CG softwares.  [47, 32, 44, 4, 92, 37] could produce animatable avatars, but they ignored garment-human interactions, leading to unsatisfactory animation quality.
Disentangled representations. Reconstructing human body and clothes as separate geometries has long been studied in the field of computer vision [18, 39, 25, 26], but generating distinct body and clothes representations has barely been explored. Previous works have made attempts to generate 3D avatars as a combination of separate components [31] or layers [86], but they did not explicitly disentangle human body and accessories, making an individual component or layer not physically meaningful. A more recent work [33] trained a diffusion-based human generation model and distinctly-represented human body and different types of clothes as separate layers, but it did not have semantic controllability and suffers from over-smoothing textures. A concurrent work TECA [93] explored disentangling hair and head ornament from text and images while our work focuses on decomposing the avatar into a human body and clothes. Another recent work AvatarFusion [34] also separates clothes from human bodies. However, AvatarFusion adopts a post-processing strategy based on the outputs of AvatarClip [32], leading to low-quality results. Our method is natively designed for disentangled generation, which produces much better quality.
Figure 2: Illustration of our SO-SMPL representation. Two vertex-wise offsets, namely the human offset $\mathbf{O}_{\text{h}}$ and the clothes offset $\mathbf{O}_{\text{c}}$ are sequentially added in order to the SMPL-X body mesh $\mathbf{T}(\beta)$ to obtain the human body mesh $\mathbf{T}_{\text{h}}(\beta,\mathbf{O}_{\text{h}})$ and clothed human mesh $\mathbf{T}_{\text{c+h}}(\beta,\mathbf{O}_{\text{h}},\mathbf{O}_{\text{c}},\mathbf{M}_{\text{c}})$, where a vertex mask $\mathbf{M}_{\text{c}}$ is calculated to determine the clothing region. Finally, we mask the clothed human mesh with $\mathbf{M}_{\text{c}}$ and obtained the separated clothes mesh $\mathbf{T}_{\text{c}}(\beta,\mathbf{O}_{\text{h}},\mathbf{O}_{\text{c}},\mathbf{M}_{\text{c}})$.
3 Preliminaries
Score Distillation Sampling [62] is proposed to optimize a 3D representation through a frozen pre-trained large 2D diffusion model. Specifically, to optimize a given 3D representation with parameters $\theta$ using a diffusion model with parameters $\phi$, SDS loss is computed as:
$\nabla_{\theta}\mathcal{L}_{SDS}=E_{t,\epsilon}\left[(\hat{\epsilon}_{\phi}(x_{t};\mathbf{y},t)-\epsilon)\frac{\partial x}{\partial\theta}\right]$ (1)
where $t$ is the time step in diffusion model, $x=g(\theta)$ is the rendered image using a differentiable renderer $g$, $x_{t}=x+\epsilon$ is the noise version of $x$, $\epsilon$ is a sampled gaussian noise, and $\hat{\epsilon}_{\phi}(x_{t};\mathbf{y},t)$ is the predicted noise of the diffusion model.
SMPL-X [60] is an expressive parametric human model that could produce a human mesh with a fixed topology. SMPL-X is conditioned by a given input shape parameter $\beta$, pose parameter $\theta$, and expression parameters $\psi$, and follows a vertex-based linear blend skinning(LBS) based on skeleton $J(\beta)$ and skinning weights $W$ for pose transforming and animation. Specifically, SMPL-X computes a posed human mesh $\mathbf{K}(\beta,\theta,\psi)$ as:
$\begin{split}\mathbf{K}(\beta,\theta,\psi)&=\mathcal{W}(\mathbf{T}(\beta,\theta,\psi),J(\beta),\theta,W)\\ \mathbf{T}(\beta,\theta,\psi)&=T+B_{s}(\beta)+B_{e}(\psi)+B_{p}(\theta)\end{split}$ (2)
where an T-pose body mesh $\mathbf{T}(\beta,\theta,\psi)$ is first calculated as the combination of a mean template $T$, shape, expression and pose blend shapes $\{B_{s},B_{e},B_{p}\}$ then warped to target pose $\theta$ with LBS operation $\mathcal{W}$.
4 Methodology
Figure 3: An overview of our generation pipeline. Our pipeline has two stages. In Stage I, we generate a base human body model by optimizing its shape parameter $\beta$, vertex offset $\mathbf{O}_{\text{h}}$ and albedo texture $\Gamma_{\text{h}}$. A ControlNet is utilized to compute a Score Distillation Sampling(SDS) Loss conditioned on an ”A”-pose map $\mathcal{P}$ for the rendered RGB image $\mathbf{I}_{\text{h}}$ and normal map $\mathbf{N}_{\text{h}}$. In Stage II, we freeze the human body model and optimize the clothes parameters $\mathbf{M}_{c}$ and $\mathbf{O}_{\text{c}}$ along with the albedo texture MLP $\Gamma_{\text{c}}$. The rendered RGB images and normal maps of both the clothed human and the clothes are used in computing the SDS losses. In both stages, we utilized a simple Phong shading model to render images from our SO-SMPL representations.
Fig. 3 is an overview of our disentangled avatar generation pipeline. Given text descriptions of the human avatar, our pipeline produces an animatable and high-quality unclothed avatar in the first stage. In the second stage, we create clothes on the target avatar. In the following, we introduce our disentangled representation, called SO-SMPL, of the human body and the clothes.
4.1 SO-SMPL representation
Human Body Geometry. We represent the human body mesh with a densified SMPL-X parametric model introduced in Eq. 2 and we add learnable vertex-wise offsets $\mathbf{O}_{\text{h}}$ to represent details on the body, similar to recent works [47]. We fix the pose of the human body mesh to an “A” pose $\theta$ and use a default expression parameter $\psi$ during the text-to-avatar generation process so that $\theta$ and $\psi$ are omitted for simplicity. Specifically, the human body geometry $\mathbf{T}_{\text{h}}$ is
$\begin{split}\mathbf{T}_{\text{h}}(\beta,\mathbf{O}_{\text{h}})&=\mathbf{T}(\beta)+\mathbf{O}_{\text{h}},\end{split}$ (3)
where $\mathbf{O}_{\text{h}}$ is trainable vertex-wise offsets and $\beta$ is the trainable human shape parameter.
Clothes Geometry. Clothes are represented by additional vertex-wise masks $\mathbf{M}_{\text{c}}$ and offsets $\mathbf{O}_{\text{c}}$ on $\mathbf{T}_{\text{h}}$. The clothed human mesh $\mathbf{T}_{\text{c+h}}$ are
$\begin{split}\mathbf{T}_{\text{c+h}}(\beta,\mathbf{O}_{\text{h}},\mathbf{O}_{\text{c}},\mathbf{M}_{\text{c}})&=\mathbf{T}_{\text{h}}(\beta,\mathbf{O}_{\text{h}})+\mathbf{O}_{\text{c}}\odot\mathbf{M}_{\text{c}},\end{split}$ (4)
where $\odot$ denotes the Hadamard product between the vertex-wise offsets and masks. We could also get the clothes geometry $\mathbf{T}_{\text{c}}$ only by masking on the clothed human mesh
$\begin{split}\mathbf{T}_{\text{c}}(\beta,\mathbf{O}_{\text{h}},\mathbf{O}_{\text{c}},\mathbf{M}_{\text{c}})&=\mathbf{T}_{\text{c+h}}(\beta,\mathbf{O}_{\text{h}},\mathbf{O}_{\text{c}},\mathbf{M}_{\text{c}})\odot\mathbf{M}_{\text{c}}.\end{split}$ (5)
To enable the geneartion of diverse clothes types, we use the body-part segments of SMPL-X and pre-define vertex-level mask templates of 6 different garment types, namely long shirt, short shirt, long pants, short pants, vest, and overalls. These templates are used to initialize clothes mask $\mathbf{M}_{\text{c}}$.
Discussion. Previous works [93, 34] tried to represent clothes as a volume-based representation, e.g. NeRF [56], which is hard to be converted to a mesh. Our clothes representation is naturally a thin mesh representation, which is preferable for most graphics software and physical simulators [2, 1]. Meanwhile, by representing the clothes as a displacement surface that “grows” upon the base human, we ensure the alignment between the human body and clothes.
4.2 Rendering SO-SMPL
On both the human body mesh and the clothes mesh, we model their appearances with a simple Phong shading model where the meshes are characterized by their albedo and the lighting is simply a combination of a point light and an ambient light.
Albedo We use an MLP $\Gamma_{\text{i}}$ to represent the albedo of the mesh. The albedo $\mathbf{\rho}$ of a 3D point $\mathbf{x}$ on the target mesh could be computed as i.e.:
$\mathbf{\rho}$ $=\Gamma_{i}(\gamma(\mathbf{x});\xi_{i})$ (6)
where $i\in\{\text{h},\text{c}\}$ means clothes or human body, $\gamma$ is positional encoding of the 3D position $\mathbf{x}$, and $\xi_{i}$ is the parameter of the MLP.
Shading model. Given a randomly-sampled point light with a position of $\mathbf{v}$ and an intensity of $\mathbf{l}_{\text{d}}$, along with the ambient light intensity $\mathbf{l}_{\text{a}}$, the shading color $\mathbf{c}$ of a surface point $\mathbf{p}$ could computed by
$\begin{split}\mathbf{c}&=\mathbf{\rho}_{\mathbf{p}}(\mathbf{l}_{\text{a}}+\max(0,\vec{\mathbf{n}}_{l}\cdot\vec{\mathbf{n}}_{\mathbf{p}})\mathbf{l}_{\text{d}})\end{split}$ (7)
where $\vec{\mathbf{n}}_{\mathbf{p}}$ is the normal direction of the point $\mathbf{p}$, and $\vec{\mathbf{n}}_{l}=(\mathbf{v}-\mathbf{p})/\|\mathbf{v}-\mathbf{p}\|_{2}$ is the light direction towards the surface vertice.
Figure 4: An intuitive illustration of the impact of the baked-in artifacts in learned albedo. On the left, the generated pant suffers from severe baked-in wrinkles in its texture, resulting in non-photorealistic wrinkles and shadows in animation. On the right side, our shader prevents the shadows from baking into the albedo, hence significantly improving the visual quality of animations.
Discussion. This shading model helps us to avoid the shadows being baked into the albedo of the generated clothes and improves the consistency between geometry and textures, as shown in Fig. 4. This enhancement can be attributed to our SO-SMPL’s ability to compute normals with greater accuracy compared to previous methods. These precisely estimated normals facilitate our shading model’s capacity to explicitly render shadows, particularly those resulting from wrinkles on the clothes.
Rendering Clothes & Human Body. From any given camera viewpoint, we can render the images from our SO-SMPL representation using rasterization. We first determine the corresponding 3D point and its normal from $\mathbf{T}_{\text{h}+\text{c}}$ on every pixel. Then, the shading colors are computed using Eq. 7, resulting in the rendered image $\mathbf{I}_{\text{h}+\text{c}}$. In a similar fashion, we render separate images for the body ($\mathbf{I}{\text{h}}$) and the clothes ($\mathbf{I}{\text{c}}$), respectively. Concurrently, we also obtain the corresponding normal maps $\mathbf{N}_{\text{h}+\text{c}}$, $\mathbf{N}_{\text{h}}$, and $\mathbf{N}_{\text{c}}$. More details are included in the supplementary material.
4.3 SO-SMPL Generation
To learn the geometry and texture representations of the SO-SMPL model, we utilize SDS losses on both the rendered RGB images and the normal maps. Our approach involves a two-stage optimization process: In Stage I, we focus on generating an unclothed human body model. Following this, in Stage II, we proceed to generate the clothing, which is attached to the human body model established in Stage I.
Stage I: Human body generation. For the human body generation, we adopt an OpenPose [10]-based ControlNet [97] with SDS [62] losses to optimize a SO-SMPL representation. An “A”-pose joint map $\mathcal{P}$ rendered from 3D skeleton joints is utilized as the condition for ControlNet to compute SDS losses. Given an input rendered image $\mathbf{I}\in\{\mathbf{I}_{\text{h}},\mathbf{N}_{\text{h}}\}$, the gradient is computed by
$\nabla_{\Theta_{\text{h}}}\mathcal{L}=E_{t,\epsilon}\left[(\hat{\epsilon}_{\phi}(\mathbf{I}_{t};\mathbf{y}_{\text{h}},\mathcal{P},t)-\epsilon)\frac{\partial\mathbf{I}}{\partial\Theta_{\text{h}}}\right],$ (8)
where $\mathbf{I}_{t}=\mathbf{I}+\epsilon$ is a noisy version of the input image, $\Theta_{\text{h}}=\{\beta,\mathbf{O}_{\text{h}},\xi_{\text{h}}\}$ is the trainable parameters during human body generation, $\{\beta,\mathbf{O}_{\text{h}}\}$ are geometry parameters, $\xi_{\text{h}}$ is the texture parameters, $\hat{\epsilon}_{\phi}(\mathbf{I}_{t};\mathbf{y}_{\text{h}},\mathcal{P},t)$ is the predicted noise given text embedding $\mathbf{y}_{\text{h}}$, noise step $t$ and the pose map $\mathcal{P}$. To ensure our generated human body mesh is unclothed, we added specific descriptions to the human prompt such as “wearing tight shorts”. Additionally, to further guide the body mesh and texture away from depicting clothing, we employed negative prompts like “loose clothes, accessories”.
Stage II: Clothes generation. Similar to human body generation, we also utilize SDS losses to update clothing parameters $\Theta_{\text{c}}=\{\mathbf{O}_{\text{c}},\xi_{\text{c}}\}$. We calculate SDS gradients through RGB and normal images of both the clothed human and the separated clothes. Given a rendered image $\mathbf{I}\in\{\mathbf{I}_{\text{c}},\mathbf{N}_{\text{c}},\mathbf{I}_{\text{c+h}},\mathbf{N}_{\text{c+h}}\}$, and its corresponding text prompt embedding $\mathbf{y}_{\text{c}}$, the SDS loss is termed as:
$\begin{split}\nabla_{\Theta_{\text{c}}}\mathcal{L}=E_{t,\epsilon}\left[(\hat{\epsilon}_{\phi}(\mathbf{I}_{t};\mathbf{y}_{\text{c}},t)-\epsilon)\frac{\partial\mathbf{I}}{\partial\Theta_{\text{c}}}\right]\end{split}$
Albedo smoothness constraint. To avoid geometries (e.g. wrinkles and bumps) and lighting (shadows and highlights) being baked into the albedo of the model, we employed an albedo smoothness loss following [98, 17] to reduce these effects. With a small perturbation $\delta$ of the input position $\mathbf{x}$, we encourage the estimated albedo from texture MLP $\Gamma_{i}\in\{\Gamma_{\text{h}},\Gamma_{\text{c}}\}$ to be consistent, i.e.
$\mathcal{L}_{a}=\|\Gamma_{i}(\mathbf{x})-\Gamma_{i}(\mathbf{x}+\delta)\|_{2}$ (9)
Geometry constraints. In both the human body and clothing generation, we utilized geometry constraints to ensure that the mesh is smooth and does not deviate too far from the original SMPL-X mesh. We employed a Laplacian smoothness term [41] $\mathcal{L}_{s}$ to minimize the norm of graph Laplacian. We also added an offset regularizing term $\mathcal{L}_{o}=\|\mathbf{O}_{i}\|_{2}$ for the body/clothes offset $\mathbf{O}_{i}\in\{\mathbf{O}_{\text{h}},\mathbf{O}_{\text{c}}\}$. Finally the geometry regularization los could be written as $\mathcal{L}_{\text{geo}}=\mathcal{L}_{s}+\mathcal{L}_{o}$.
Figure 5: Qualitative comparisons of animation results. Our model enables the generation of disentangled human body and clothes meshes, which thus enables more photorealistic animations.
5 Experiments
5.1 Generated Avatars & Clothes
Figure 6: Our method is able to generate avatars with varied ethnicities, genders, and clothes.
As demonstrated in Fig. 6, our method generates diverse avatars and various types of clothes with detailed texture. Besides, since the clothes were generated on a base human body, they fit with the intended human naturally.
5.2 Quality Comparisons
We conducted comparisons on three aspects, static clothed-avatar quality, clothes quality, and animation quality. For the generation quality of clothed avatars, we compared with state-of-the-art text-to-3D avatar generation methods TEXture [66], AvatarCLIP [32] and TADA! [47]. As for clothes generation, we conducted comparisons with two of the best existing text-to-3D methods ProlificDreamer [82] and TextMesh [75]. We then compared animation results in the physical simulation environment [2] with the animatable meshes generated by TEXture [66] and TADA! [47].
Clothed Avatar Quality.
Figure 7: Quality comparison of the generated clothed avatars from TEXture [66], AvatarCLIP [32], TADA! [47] and our method with the same prompts.
A visual comparison of the generated avatars is presented in Fig. 7. TEXTure [66] takes a fixed SMPL-X mesh as input and suffers from texture inconsistency under different views. AvatarCLIP [32] often produces geometry artifacts and poor facial details. TADA! [47] tends to generate unnatural geometric artifacts and unaligned outfits, e.g. a hull on the African-American woman’s belly and the white teenage girl’s yellow pants that contradicts the “beige legging” described by the prompt. In comparison, our method is able to produce not only the most realistic human with detailed textures but also high-quality clothes that are semantically aligned with texts.
Clothes Quality. We also evaluate the quality of the separated clothes. We compared with state-of-the-art text-to-3D generation methods ProlificDreamer [82] and TextMesh [75] for standalone clothes generation. As demonstrated in Fig. 8, neither TextMesh [75] or ProlificDreamer [82] could generate meaningful clothes shapes & textures. Besides, they also could not control the position and the size of the generated clothes, and thus could not be fitted to an avatar. Meanwhile, our method is able to produce clothes fitting to a specific avatar ana also with high-quality textures and shapes.
Figure 8: Comparisons of clothes generation quality using the same prompts.
Animation Quality. Our generated human body mesh is built upon the SMPL-X model and thus allows us to use SMPL-X parameters to animate the body. On the other hand, the separately generated clothes can be used in a simulation environment [2, 1] to simulate their motions. We used motion sequences from AMASS [54] and AIST++ [46] to animate the generated avatars and simulate the clothes animations in the MarvelousDesigner [2]. Note that the recent progress in human motion synthesis [22, 74, 76, 61, 100, 3] or human motion capture from visual signals [45, 15, 94, 23, 81, 71] can also be employed to drive the generated character. The animation results of TADA!, TEXTure, and our method are shown in Fig. 5. For TADA! and TEXTure, the clothes are entangled on the generated avatar, which leads to inconsistent wrinkles and shadows in animation. Our method disentangles clothes and human bodies, enabling realistic clothes-body interactions and visual effects.
5.3 User Study
A comprehensive user study is conducted to evaluate the generation quality of our pipeline. We designed a questionnaire and presented 16 full-body animation made by TEXture [66], TADA! [47], and our method. The survey was conducted among 26 researchers and artists from either academia or industry. The participants are asked to pick the one with the best semantic consistency, static visual quality, and animation quality. As shown in Tab 1, our model was the most preferred one in all aspects of generation quality.
Favourite (%, $\uparrow$) TEXture [66] TADA [47] Ours
Texture & Geometry Quality 8.1 8.4 83.4
Animation Realism 3.4 7.0 89.7
Prompt Alignment 5.5 7.7 86.8
Table 1: User study from 26 users in terms of visual quality, animation quality, and prompt alignment preference. The number means the percentage of users selecting a specific method.
5.4 Ablation Study
We evaluated the effects of our albedo smoothness constraints and 2D-pose ControlNet guidance. In Fig. 9, we compared the rendered albedo and RGB images w/ or w/o our albedo smoothness constraint $\mathcal{L}_{a}$. As can be seen from the visual results, the albedo generated without $\mathcal{L}_{a}$ inherent shadows and wrinkles, which would lead to unsatisfactory visual artifacts in animations. In comparison, using $\mathcal{L}_{a}$ largely alleviates the problem in the albedo while still enabling the correct rendering of the shadows and wrinkles on RGB images.
Figure 9: Ablation study on our shading model. For each garment, we showcase the albedo map (left) and the rendered RGB image (right). Our full shading model with albedo smoothness constraint $\mathcal{L}_{a}$ enables the rendering of shadows in wrinkles while alleviating baked-in artifacts of shadows in the albedo.
In Fig. 10, we validated the effectiveness of Openpose-based ControlNet as guidance during our human body generation. As could be observed from the results, generated avatars without ControlNet have oversmoothed textures and geometry. In contrast, our utilization of ControlNet produces significantly finer details in both geometry and texture.
Figure 10: Ablation study on ControlNet as guidance. We compare the avatars generated by the ControlNet with the ones generated by the vanilla Stable diffusion model. The results showed that the ControlNet model with 2D pose conditions generates avatars with more detailed geometry and texture.
5.5 Applications
Virtual Try-on. Our disentangled human-clothes representation inherently allows us to change the outfit of a certain avatar, or put the same clothes on different avatars.
Figure 11: Comparisons in virtual try-on applications. The orange and purple arrows represent compositions of the human body and the clothes, respectively. Changing clothes in TADA! [47] often leads to undesirable changes in human identity or clothes while changing clothes in our method is more photorealistic and keeps the same human identity and clothes.
As depicted in Fig. 11, while previous methods like TADA! [47] also enable virtual try-on, they struggle to maintain consistency in clothing and the human identity during the try-on process. As can be seen from the left half of Fig. 11, when changing clothes for the white teenage boy with TADA!, the facial details and body shape are changed; on the other hand, the colors and patterns of the clothes also changed when putting the same clothes on different avatars with TADA!. On the contrary, our disentangled representation supports fixing the human identity and clothes details when switching outfits.
Clothes Shape Editing Our offset-based model also allows the same outfit design to be fitted on people of different shapes. Under our SO-SMPL representation, we could simply combine the human body mesh parameters of the target avatar: the SMPL-X shape $\beta$ and offset $\mathbf{O}_{\text{h}}$, with the geometry offset $\mathbf{O}_{\text{c}}$ and texture MLP $\Gamma_{\text{c}}$ of the intended clothes. As shown in Fig. 12, clothes generated by our methods could be changed according to different human shapes and sizes without re-generating. This flexibility helps our generated clothes to be re-used on different avatars while maintaining fit with their body types.
Figure 12: Shape editing on clothes. Our SO-SMPL representation allows us to fit generated clothes onto different human shapes while maintaining the original patterns and details.
6 Limitations&Conlusion
Limitations. Despite promising results, our pipeline still has limitations. Currently, since our SO-SMPL representation is based on SMPL-X model, our pipeline could not generate clothes types that do not follow SMPL-X’s topological structure, such as skirts and dresses. To extend to such clothes, we may follow [39] and utilize a parametric dress/skirt model as template instead of original SMPL-X. Another limitation is that our generated clothes do not follow any sewing patterns, and are without any physical attributes such as stretch, bending, and friction. Therefore, generating garments with sewing patterns and physical properties is a promising direction for exploration. We leave these potential tasks to future works.
Conclusion. In this paper, we have presented an innovative approach for generating human avatars through our Sequentially Offset-SMPL (SO-SMPL) representation from textual descriptions. To the best of our knowledge, our pipeline is the first to produce avatars in a disentangled manner: first generate a human body mesh then a clothes mesh on top of it. The 3D avatars and clothes generated by our pipeline exhibit remarkable diversity and high fidelity in both texture and geometric detailing. They could also be easily utilized in CG software for animation and simulation, thereby opening up exciting possibilities in fields such as virtual reality, gaming, and digital fashion. Experiments demonstrate that our pipeline outperforms existing text-to-3D generation methods in texture&geometry quality, alignment with text descriptions, and animation quality.
Supplementary Material
7 Pipeline Details
7.1 Mask Templates
As mentioned in Sec. 4.1, we utilized SMPL-X[53] part segments to craft templates for 6 different types of clothes. The types of masks could be seen from Fig. 13.
Figure 13: Mask templates.
7.2 Negative Prompting
Simply using descriptive prompts can not guarantee the distinction between human body and clothes. For instance, even with descriptions like “unclothed” or “without accessories”, there could still be undesired extras on the human body. On the other hand for garment generation, the generated clothes sometimes contain undesired “human parts” like skin or even limbs. To address this problem, we added negative prompts for both body(e.g. “loose clothes, accessories”) and clothes generation(e.g. “skin, legs”), to encourage the predicted noise from Stable Diffusion to guide the model away from such entanglement. Specifically, the original classfier-free guidance[29] against unconditoned noise $\epsilon_{\phi}(\mathbf{I}_{t};t)$ is:
$\hat{\epsilon}_{\phi}(\mathbf{I}_{t};\mathbf{y},t)=(1+\omega)\epsilon_{\phi}(\mathbf{I}_{t};\mathbf{y},t)-\omega\epsilon_{\phi}(\mathbf{I}_{t};t)$ (10)
and the modified version with negative prompting is:
$\hat{\epsilon}_{\phi}(\mathbf{I}_{t};\mathbf{y},t)=(1+\omega)\epsilon_{\phi}(\mathbf{I}_{t};\mathbf{y},t)-\omega\epsilon_{\phi}(\mathbf{I}_{t};\mathbf{y}_{\text{n}},t)$ (11)
7.3 Mask Blending
To render the clothed human rgb image $\mathbf{I}_{\text{c+h}}$ from separate body and clothes textures, we rasterize the SO-SMPL with its vertex-level mask $\mathbf{M}_{\text{c}}$ to obtain 2D blending mask $\mathbf{I}^{\text{mask}}$ and use it to combine the rgb values of human body $\mathbf{I}_{\text{h}}$ clothing $\mathbf{I}_{\text{c}}$, i.e. :
$\begin{split}\mathbf{I}_{\text{c+h}}&=\mathbf{I}_{\text{c}}\odot\mathbf{I}^{\text{mask}}+\mathbf{I}_{\text{h}}\odot(\mathbf{1}-\mathbf{I}^{\text{mask}})\end{split}$ (12)
where $\odot$ denotes the Hardmard product. Notably, self-occlusions are already considered in the rasterization process thus $\mathbf{I}^{\text{mask}}$ could be applied directly.
8 Additional Experiments
Figure 14: Generated human body meshes.
Figure 15: An overall view of our generated clothes and avatars.
Figure 16: Generated clothes gallery.
8.1 Experimental Details
All experiments for this paper are conducted on a single GeForce RTX 4090 GPU. For each human body mesh, it takes around 12GB GPU memory and 2.5 hours to generate. As for clothes, it takes around 19GB GPU memory and 2.5 hours of optimizing.
8.2 Avatar Gallery
A gallery of our generated avatars is presented in Fig. 14. Our method is able to generate highly detailed human body meshes.
8.3 Clothes Gallery
A gallery of our generated clothes of different types is presented in Fig. 16. As could be seen from the results, our pipeline could generate a diverse range of clothes of different colors, materials, and types.
8.4 Additional Qualitative Comparisons with DreamWaltz [37]
We present comparisons with an additional state-of-the-art text-to-avatar method, namely DreamWaltz [37]. Visual comparisons are presented in Fig. 18 and Fig. 17.
As seen in Fig. 17, DreamWaltz [37] is capable of generating photorealistic plausible results, yet it suffers from over-smooth texture and noises due to its NeRF-based representation. Consequently in Fig. 18, noises move along with the joint motions, causing severe visual artifacts.
Figure 17: Comparisons of “A-pose” quality with DreamWaltz [37].
Figure 18: Comparisons of animation quality with DreamWaltz [37]. For a better visual comparison, please refer to our supplementary videos.